OpenStudy (anonymous):

an object is thrown straight upward. It's height in feet (y) after seconds is given by y=-6t2+128t. What is the maximum height reached by the object? How long does it take to reach that height?

OpenStudy (anonymous):

easiest way is to use t=-b/(2a) then solve for y b=128 a=-6 so t=-128/(2*-6)=32/3 no solve for y y=-6(32/3)^2+128(32/3) y=(-2048/3)+(4096/3) y=(4096-2048)/3 y=2048/3